# Learning parameter dependence for Fourier-based option pricing with tensor trains

### v7

@@ V.1 Details
Ranges of ${\sigma_{m}}$ and ${S_{m,0}}$
@@
Other parameters
The other parameters for option pricing are fixed to the values summarized in the table below.
$T$ $r$ $K$ $S_{0}$ $\alpha$ $\rho_{mn},m\neq n$ $(N,\eta)$
$1$ $0.01$ $100$ $100$ $\frac{5}{d}$ $\frac{1}{3}$$\left\{\begin{array}[]{ll}(200,0.3)&;\ d=11,\vec{p}=\vec{S}_{0}\\ (100,0.4)&;\ \mathrm{otherwise}\end{array}\right.$
Table 1: The parameters except $\vec{\sigma}$ and $\vec{S}_{0}$.
Error evaluation
@@
Software and hardware used in this study
TensorCrossInterpolation.jl [36] was used for learning tensor trains from functions appearing Fourier-based option pricing. ITensors.jl [37] was utilized for computing the contraction and inner product of the tensor trains. The Monte Carlo simulations were carried out using tf-quant-finance [38]. Parallelization was not employed in either case. GPUs were not utilized in any of the calculations. The computations were performed on a 2023 MacBook Pro featuring a 12-core Apple M2 Max processor and 32GB of 400GB/s unified memory.
VI Results
@@
The results are summarized in TABLE 2. In particular, the computational complexity versus $d$ is plotted in FIG. 2. The mean absolute error in TT-based option prices among runs for 100 random parameter sets is represented by $e_{\mathrm{TT}}$. For the same parameter sets, we compute mean values of the absolute values of the error in the MC-based option price with $10^{6}$ paths and denote it by $e_{\mathrm{MC}}$. The computational complexity and time of TT-based option pricing are represented by $c_{\mathrm{TT}}$ and $t_{\mathrm{TT}}$, respectively, and those of MC-based pricing with $10^{6}$ paths are denoted by $c_{\mathrm{MC}}$ and $t_{\mathrm{MC}}$. To maintain the desired accuracy of option pricing, we set the tolerance of TCI sufficiently low, concretely $\epsilon_{\mathrm{TCI}}=10^{-9}$. Subsequently, we reduce the bond dimension by SVD, with the tolerance of SVD set to $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}=1.0\times 10^{-6}$ for $\phi$ and $\hat{v}$. The maximum bond dimensions of the TTs for $\phi$ and $\hat{v}$ are denoted as $\chi_{\phi}$ and $\chi_{\hat{v}}$, respectively, in TABLE 2. In the SVD on $\tilde{V}_{k_{1},\ldots,k_{d}}$ obtained by contracting the TTs for $\phi$ and $\hat{v}$, we set the tolerance $\epsilon^{\tilde{V}}_{\mathrm{SVD}}=1.0\times 10^{-6}$, and the maximum bond dimension of the resulting TT is denoted by $\chi_{V}$.
VI.0.1 The case of varying $\vec{\sigma}$
TABLE 2 (a) shows the computational results of TT-based option pricing when we consider $\vec{\sigma}$ dependence. TT-based option pricing demonstrates advantages in terms of computational complexity and time over the MC-based method. The bond dimensions of the TT results for $d=5$ and $d=10$ are depicted in FIG. 3(a). By applying SVD to the TTs trained via TCI, the bond dimensions between tensors related to $p_{m}$ and $z_{m+1}$ could generally be maintained at around 10 when $\eta$ is constant. The details of this compression by SVD are described in Appendix. A. The maximum bond dimension $\chi_{\tilde{V}}$ of $\tilde{V}_{k_{1},\ldots,k_{d}}$ is maintained at 2, or 1 especially when $d=10,11$. Therefore, the computational complexity is much lower than that of MC-based option pricing. Besides, the error in the TT-based result is smaller than that in the MC-based result in all the tested cases.
@@ VI.0.1 The case of varying $\vec{\sigma}$
$d$ $e_{\mathrm{TT}}$ $e_{\text{MC}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC}}$ $t_{\mathrm{TT}}$[s] $t_{\mathrm{MC}}$[s] $\chi_{\phi}$ $\chi_{\hat{v}}$ $\chi_{\tilde{V}}$
$5$ 0.00178 0.00606 $16$ $5.0\times 10^{6}$ $4.91\times 10^{-7}$ $2.84\times 10^{-1}$ 16 11 2
$6$ 0.00154 0.00503 $20$ $6.0\times 10^{6}$ $6.37\times 10^{-7}$ $3.38\times 10^{-1}$ 16 11 2
$7$ 0.00134 0.00428 $24$ $7.0\times 10^{6}$ $9.12\times 10^{-7}$ $3.71\times 10^{-1}$ 17 11 2
$8$ 0.00136 0.00372 $28$ $8.0\times 10^{6}$ $1.75\times 10^{-6}$ $4.09\times 10^{-1}$ 18 11 2
$9$ 0.000867 0.00329 $32$ $9.0\times 10^{6}$ $2.47\times 10^{-6}$ $4.51\times 10^{-1}$ 20 11 2
$10$ 0.00229 0.00294 $10$ $1.0\times 10^{7}$ $1.18\times 10^{-6}$ $5.11\times 10^{-1}$ 20 11 1
$11$ 0.000554 0.00265 $11$ $1.1\times 10^{7}$ $1.16\times 10^{-6}$ $5.04\times 10^{-1}$ 20 11 1
@@
$d$ $e_{\mathrm{TT}}$ $e_{\text{MC}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC}}$ $t_{\mathrm{TT}}$[s] $t_{\mathrm{MC}}$[s] $\chi_{\phi}$ $\chi_{\hat{v}}$ $\chi_{\tilde{V}}$
$5$ 0.00151 0.00773 $16$ $5.0\times 10^{6}$ $7.67\times 10^{-7}$ $3.99\times 10^{-1}$ 18 11 2
$6$ 0.00122 0.00640 $20$ $6.0\times 10^{6}$ $6.34\times 10^{-7}$ $5.15\times 10^{-1}$ 19 11 2
$7$ 0.00112 0.00547 $24$ $7.0\times 10^{6}$ $9.21\times 10^{-7}$ $5.37\times 10^{-1}$ $21$ $10$ $2$
$8$ 0.000973 0.00477 $28$ $8.0\times 10^{6}$ $1.63\times 10^{-6}$ $6.18\times 10^{-1}$ $23$ $11$ $2$
$9$ 0.000686 0.00424 $32$ $9.0\times 10^{6}$ $1.19\times 10^{-6}$ $7.95\times 10^{-1}$ $24$ $11$ $2$
$10$ 0.000662 0.00377 $36$ $1.0\times 10^{7}$ $1.47\times 10^{-6}$ $8.77\times 10^{-1}$ $24$ $11$ $2$
$11$ 0.00114 0.00339 $40$ $1.1\times 10^{7}$ $1.93\times 10^{-6}$ $8.43\times 10^{-1}$ $25$ $13$ $2$
Table 2: The results of TT-based option pricing incorporating (a) $\vec{\sigma}$ and (b) $\vec{S}_{0}$ dependence. Here, we set the ranges $\sigma_{m}\in[0.15,0.25)$ and $S_{m,0}\in[90,120)$, and place 100 equally spaced grid points within these ranges. $t_{\mathrm{TT}}$ and $t_{\mathrm{MC}}$ represent the average computational time from 100 measurements, respectively.
Figure 2: The computational complexity of TT-based and MC-based option pricing, denoted as $c_{\mathrm{TT}}$ and $c_{\mathrm{MC}}$ versus the number of assets $d$. We consider including parameter dependencies on $\vec{\sigma}$ and $\vec{S_{0}}$ in TTs and the number of path in MC are $10^{6}$.
@@
VI.0.2 The case of varying $\vec{S_{0}}$
@@ VI.0.3 Randomness in learning the TTs
VI.0.4 Total computational time for obtaining the TTs
We mention the total computational time for obtaining the TTs for $\phi$ and $\hat{v}$ and $\tilde{V}_{k_{1},\ldots,k_{d}}$ through TCI and SVD. We focus on TCI because it dominates over SVD. It took about $12.5$ minutes for $d=11$ in the case that $\vec{S}_{0}$ dependence is involved. This is sufficiently short for a practical use-case of the TT-based method mentioned exemplified in Sec. VII. We also note that, when we do not incorporate the parameter dependence into the tensor train for $\phi$ as in Ref. [6], TCI takes a much longer time than the MC-based method, 7.3 seconds for $d=11$ in the case of $\vec{\sigma}$ dependence. In this setup, we set $\eta=0.48$, $\epsilon_{\mathrm{TCI}}=10^{-6}$, $N=50$ and all other parameters are the same as those in the default settings listed in Table 1. This means that running TCI every time the parameter varies does not lead to the time advantage of the TT-based method over the MC-based one.
VII Discussion
@@
Finally, we discuss future research directions. For our method to be more practical, compressing all the input parameters including both $\vec{\sigma}$ and $\vec{S}_{0}$ into a single TN format is desired, although this study tried to incorporate either of them as the first step. Besides, to enhance the practical benefit, we should extend our methodology so that it is applicable to more advanced settings in option pricing. With respect to the pricing model, while this paper has considered the BS model, more advanced models such as the local volatility model [39], stochastic volatility model [40, 41] and Lévy models [42, 43] are also used in practice. Applying our method to such models, where the dependencies on the parameters in the models are incorporated into TTs, is an interesting direction. Expanding the scope to the broad types of products is also desired. While this paper focused on European-type options, more complicated ones such as American and Bermudan options are also traded widely, and pricing of such options is more time-consuming. Considering whether our method can be extended for such options will be an important and interesting challenge.
Acknowledgements
@@ Appendix A How to set the tolerance $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$
Figure 4: The mean absolute error $e_{\mathrm{TT}}$ when applying SVD with various $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ values to the TTs when incorporating the $\vec{\sigma}$ dependence at $d=10$.
Figure 4 shows the maximum relative error $e_{\mathrm{TT}}$ when applying SVD with varying $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ to the TT when incorporating $\vec{\sigma}$ dependence at $d=10$ obtained through TCI. Here, we fixed the tolerance of SVD $\epsilon^{\tilde{V}}_{\mathrm{SVD}}$ to $10^{-6}$. The tolerance $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ for SVD was chosen to maintain $e_{\mathrm{TT}}$ smaller than $e_{\mathrm{MC}}$. From FIG. 4, it can be seen that $e_{\mathrm{TT}}$ increases sharply between $\epsilon_{\mathrm{SVD}}=10^{-6}$ and $10^{-2}$, suggesting that setting $\epsilon_{\text{SVD}}$ around $1.0\times 10^{-6}$ is appropriate for keeping $e_{\mathrm{TT}}$ smaller than $e_{\mathrm{MC}}$. By compressing the bond dimension with SVD, the computational time for the contraction of these optimized TTs also can be kept low.
From the fact that we can keep the error $e_{\mathrm{TT}}$ small through SVD, it is suggested that the TTs obtained by TCI contain redundant information. By using SVD to get an optimal approximation in terms of the Frobenius norm, the reduced redundant information could be effectively removed. In addition, it is surprising that in the analysis with $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ = $10^{-6}$, the $e_{\mathrm{TT}}$ decreased compared to before compression by SVD. We consider that the error contained in TTs obtained by TCI is eliminated through SVD by chance. We expect that this phenomenon does not occur generally, and in fact, it did not occur for other asset numbers or parameters.

### v8

@@ V.1 Details
V.1.1 Ranges of volatility and initial stock price
@@
V.1.2 Other parameters
The other parameters for option pricing are fixed to the values summarized in the table below. The values of $\alpha$ are adopted from previous study [11] and we confirmed that the solution remains stable when the values are varied slightly. $N$ and $\eta$ are set sufficiently large and small, respectively, so that the accuracy of the proposed approach becomes better than the MC.
$T$ $r$ $K$ $S_{0}$ $\sigma$ $\alpha$ $\rho_{mn}\ (m\neq n)$ $(N,\eta)$
$1$ $0.01$ $100$ $100$ $0.2$ $\frac{5}{d}$ $\frac{1}{3}$$\begin{cases}(200,0.3)&\text{if }d=11,\ \vec{p}=\vec{S}_{0}\\ (100,0.4)&\text{otherwise}\end{cases}$
Table 1: The input parameters used in the experiment, except for $\vec{\sigma}$ and $\vec{S}_{0}$. When either $\vec{\sigma}$ or $\vec{S}_{0}$ is selected as the parameter dependency $\vec{p}$, the remaining one is fixed at the value shown in the table. For the case where $\vec{S}_{0}$ is varied and $d=11$, we use $(N,\eta)=(200,0.3)$. In all other cases, we use $(N,\eta)=(100,0.4)$.
V.1.3 Error evaluation
@@
V.1.4 Software and hardware used in this study
TensorCrossInterpolation.jl [38] was used for learning tensor trains from functions appearing Fourier-based option pricing. The Monte Carlo simulations were carried out using tf-quant-finance [39]. Parallelization was not employed in either case. GPUs were not utilized in any of the calculations. The computations were performed on a 2023 MacBook Pro featuring a 12-core Apple M2 Max processor and 32GB of 400GB/s unified memory.
V.2 Results
@@
$d$ $e_{\mathrm{TT}}$ $e_{\mathrm{MC}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC}}$ $t_{\mathrm{TT}}\,[$s]]]] $t_{\mathrm{MC}}\,[$s]]]] $\chi_{\phi}$ $\chi_{\hat{v}}$ $\chi_{\tilde{V}}$
5 0.00178 0.00606 16 $5.0\times 10^{6}$ $4.91\times 10^{-7}$ $2.84\times 10^{-1}$ 16 11 2
6 0.00154 0.00503 20 $6.0\times 10^{6}$ $6.37\times 10^{-7}$ $3.38\times 10^{-1}$ 16 11 2
7 0.00134 0.00428 24 $7.0\times 10^{6}$ $9.12\times 10^{-7}$ $3.71\times 10^{-1}$ 17 11 2
8 0.00136 0.00372 28 $8.0\times 10^{6}$ $1.75\times 10^{-6}$ $4.09\times 10^{-1}$ 18 11 2
9 0.000867 0.00329 32 $9.0\times 10^{6}$ $2.47\times 10^{-6}$ $4.51\times 10^{-1}$ 20 11 2
10 0.00229 0.00294 10 $1.0\times 10^{7}$ $1.18\times 10^{-6}$ $5.11\times 10^{-1}$ 20 11 1
11 0.000554 0.00265 11 $1.1\times 10^{7}$ $1.16\times 10^{-6}$ $5.04\times 10^{-1}$ 20 11 1
@@
$d$ $e_{\mathrm{TT}}$ $e_{\mathrm{MC}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC}}$ $t_{\mathrm{TT}}\,[$s]]]] $t_{\mathrm{MC}}\,[$s]]]] $\chi_{\phi}$ $\chi_{\hat{v}}$ $\chi_{\tilde{V}}$
5 0.00151 0.00773 16 $5.0\times 10^{6}$ $7.67\times 10^{-7}$ $3.99\times 10^{-1}$ 18 11 2
6 0.00122 0.00640 20 $6.0\times 10^{6}$ $6.34\times 10^{-7}$ $5.15\times 10^{-1}$ 19 11 2
7 0.00112 0.00547 24 $7.0\times 10^{6}$ $9.21\times 10^{-7}$ $5.37\times 10^{-1}$ 21 10 2
8 0.000973 0.00477 28 $8.0\times 10^{6}$ $1.63\times 10^{-6}$ $6.18\times 10^{-1}$ 23 11 2
9 0.000686 0.00424 32 $9.0\times 10^{6}$ $1.19\times 10^{-6}$ $7.95\times 10^{-1}$ 24 11 2
10 0.000662 0.00377 36 $1.0\times 10^{7}$ $1.47\times 10^{-6}$ $8.77\times 10^{-1}$ 24 11 2
11 0.00114 0.00339 40 $1.1\times 10^{7}$ $1.93\times 10^{-6}$ $8.43\times 10^{-1}$ 25 13 2
Table 2: The results of TT-based option pricing incorporating (a) $\vec{\sigma}$ and (b) $\vec{S}_{0}$ dependence. Here, we set the ranges $\sigma_{m}\in[0.15,0.25)$ and $S_{m,0}\in[90,120)$, placing 100 equally spaced grid points within each range. $t_{\mathrm{TT}}$ and $t_{\mathrm{MC}}$ are averages over 100 measurements.
The results are summarized in TABLE 2. In particular, the computational complexity versus $d$ is plotted in FIG. LABEL:fig:table_cc. The mean absolute error in TT-based option prices among runs for 100 random parameter sets is represented by $e_{\mathrm{TT}}$. For the same parameter sets, we compute mean values of the absolute values of the error in the MC-based option price with $10^{6}$ paths and denote it by $e_{\mathrm{MC}}$. The computational complexity and time of TT-based option pricing are represented by $c_{\mathrm{TT}}$ and $t_{\mathrm{TT}}$, respectively, and those of MC-based pricing with $10^{6}$ paths are denoted by $c_{\mathrm{MC}}$ and $t_{\mathrm{MC}}$. To maintain the desired accuracy of option pricing, we set the tolerance of TCI sufficiently low, concretely $\epsilon_{\mathrm{TCI}}=10^{-9}$. Subsequently, we reduce the bond dimension by SVD, with the tolerance of SVD set to $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}=1.0\times 10^{-6}$ for $\phi$ and $\hat{v}$. The maximum bond dimensions of the TTs for $\phi$ and $\hat{v}$ are denoted as $\chi_{\phi}$ and $\chi_{\hat{v}}$, respectively, in TABLE 2. In the SVD on $\tilde{V}_{k_{1},\ldots,k_{d}}$ obtained by contracting the TTs for $\phi$ and $\hat{v}$, we set the tolerance $\epsilon^{\tilde{V}}_{\mathrm{SVD}}=1.0\times 10^{-6}$, and the maximum bond dimension of the resulting TT is denoted by $\chi_{V}$.
V.2.1 The case of varying volatilities
TABLE 2 (a) shows the computational results of TT-based option pricing when we consider $\vec{\sigma}$ dependence. TT-based option pricing demonstrates advantages in terms of computational complexity and time over the MC-based method. The bond dimensions of the TT results for $d=5$ and $d=10$ are depicted in FIG. 2(a). By applying SVD to the TTs trained via TCI, the bond dimensions between tensors related to $p_{m}$ and $z_{m+1}$ could generally be maintained at around 10. The details of this compression by SVD are described in A. The maximum bond dimension $\chi_{\tilde{V}}$ of $\tilde{V}_{k_{1},\ldots,k_{d}}$ is maintained at 2, or 1 especially when $d=10,11$. Therefore, the computational complexity is much lower than that of MC-based option pricing. Besides, the error in the TT-based result is smaller than that in the MC-based result in all the tested cases.
@@ V.2.1 The case of varying volatilities
V.2.2 The case of varying initial stock prices
@@ V.2.3 Randomness in learning the TTs
V.2.4 Total computational time for obtaining the TTs
We mention the total computational time for obtaining the TTs for $\phi$ and $\hat{v}$ and $\tilde{V}_{k_{1},\ldots,k_{d}}$ through TCI and SVD. We focus on TCI because it dominates over SVD. It took about $12.5$ minutes for $d=11$ in the case that $\vec{S}_{0}$ dependence is involved. This is sufficiently short for a practical use-case of the TT-based method mentioned exemplified in Discussion. We also note that, when we do not incorporate the parameter dependence into the tensor train for $\phi$ as in Ref. [11], TCI takes a much longer time than the MC-based method, 7.3 seconds for $d=11$ in the case of $\vec{\sigma}$ dependence. In this setup, we set $\eta=0.48$, $\epsilon_{\mathrm{TCI}}=10^{-6}$, $N=50$ and all other parameters are the same as those in the default settings listed in Table 1. This means that running TCI every time the parameter varies does not lead to the time advantage of the TT-based method over the MC-based one.
VI Discussion
@@
Finally, we discuss future research directions. For our method to be more practical, compressing many kinds of input parameters including both $\vec{\sigma}$ and $\vec{S}_{0}$ into a single TT is desired. For example, taking into account the dependence on the parameters concerning the option contract, such as the maturity $T$ and the strike $K$, enables us to price different option contracts with a single set of TTs. Although this is a promising approach, there might be some issues. For example, it is non-trivial whether TTs incorporating many parameter dependencies have a low-rank structure. Thus, we will leave such a study for future work. Besides, to enhance the practical benefit, we should extend our methodology so that it is applicable to more advanced settings in option pricing. With respect to the pricing model, while this paper has considered the BS model, more advanced models such as the local volatility model [40], stochastic volatility model [41, 42] and Lévy models [43, 44] are also used in practice. Applying our method to such models, where the dependencies on the parameters in the models are incorporated into TTs, is an interesting direction. Expanding the scope to the broad types of products is also desired. While this paper focused on European-type options, more complicated ones such as American and Bermudan options are also traded widely, and pricing of such options is more time-consuming. Considering whether our method can be extended for such options will be an important and interesting challenge.
VII Acknowledgements
@@ Appendix A How to set the tolerance $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$
Figure 3: The maximum error $e_{\mathrm{TT}}$ when applying SVD with various $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ values to the TTs under the parameter settings of $d=10,\vec{\sigma},\eta=\text{auto}$.
Figure 3 shows the maximum relative error $e_{\mathrm{TT}}$ when applying SVD with varying $\epsilon_{\mathrm{SVD}}$ to the TT with parameters for $d=10,~{}\vec{\sigma}$, obtained through TCI. Here, we fixed the tolerance of SVD $\epsilon^{\tilde{V}}_{\mathrm{SVD}}$ to $10^{-9}$. The tolerance $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ for SVD was chosen to maintain $e_{\mathrm{TT}}$ smaller than $e_{\mathrm{MC},10^{5}}$. From FIG. 3, it can be seen that $e_{\mathrm{TT}}$ increases sharply between $\epsilon_{\mathrm{SVD}}=10^{-6}$ and $10^{-5}$, suggesting that setting $\epsilon_{\text{SVD}}$ around $5.0\times 10^{-6}$ is appropriate for keeping $e_{\mathrm{TT}}$ smaller than $e_{\mathrm{MC},10^{5}}$. By compressing the bond dimension with SVD, the computational time for the contraction of these optimized TTs also can be kept low.
From the fact that we can keep the maximum relative error $e_{\mathrm{TT}}$ small through SVD, it is suggested that the TTs obtained by TCI contain redundant information. By using SVD to get an optimal approximation in terms of the Frobenius norm, the reduced redundant information could be effectively removed. In addition, it is surprising that in the analysis with $\epsilon_{\text{SVD}}=10^{-6}$, the maximum relative error decreased compared to before compression by SVD. We consider that the error contained in TTs obtained by TCI is eliminated through SVD by chance. We expect that this phenomenon does not occur generally, and in fact, it did not occur for other asset numbers or parameters.